OpenStudy (anonymous):

draw the graph of: y=sin0{0≤0≤720}

OpenStudy (anonymous):

Dammit... Let me make it abit more... clear: If you mean something like \[ y=sin(\theta ^\circ) \quad \quad \{ 0^\circ≤ \theta ≤720^\circ \} \] Then since 720 / 180 = 4 we can say: \[ y=sin(x) \quad \quad \{ 0≤x≤4 \pi \} \] To get this in radians.. and now it looks like this: http://www.wolframalpha.com/input/?i=sin%28x%29%2C+0%3Cx%3C4pi The function of sinus, but with domain of \(0\) to \(4π\)

OpenStudy (anonymous):

what is the ampliturde and period?

OpenStudy (anonymous):

Well, you just limited the domain.. this has same period and amplitude of sin(x). So the pick amplitude is 1 to -1. pick to pick amplitude is 2.. and the period is \(2 \pi \)

OpenStudy (anonymous):

so is it 2pi sin 2?

OpenStudy (anonymous):

function: \( y = sin(x) \) amplitude: 1 to -1 pick to pick amplitude: |1| + |-1| = 2 period \(2 \pi \) Not sure what you meant.

OpenStudy (anonymous):

If they are asking for a single value of peak amplitude they probably wanna hear 1 (not -1). peak to peak amplitude is still 2 though.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Amplitude Look at the picture there, useful.

OpenStudy (anonymous):

what i mean is: